what are the odds of 1 person on a 15 member team being selected for random drug testing (10 % of team picked each Year) 4 years in a row
Answer by mathmate(429)


Question:
what are the odds of 1 person on a 15 member team being selected for random drug testing (10 % of team picked each Year) 4 years in a row

Solution:
Probability of being chosen in a particular year = 1/10.
Probability of being chosen four years in a row = (1/10)^4=1/10000
Odds of being chosen four years in a row = 1:(10000-1)=1:9999